OpenStudy (anonymous):

x^2 + 2x - 48= 0 how can you solve this

OpenStudy (michele_laino):

here, we have to apply the standard formula: \[x = \frac{{ - b \pm \sqrt {{b^2} - 4 \times a \times c} }}{{2 \times a}}\] where \(a=1, b=2,c=-48\)

OpenStudy (anonymous):

Let's solve your equation step-by-step. x2+2x−48=0 Step 1: Factor left side of equation. (x−6)(x+8)=0 Step 2: Set factors equal to 0. x−6=0 or x+8=0 x=6 or x=−8

Nnesha (nnesha):

or you can factor it find two numbers when you multiply them you should get product of AC (a=leading coefficient C =constant ) and when you add or subtract them you should get middle term ( b = coefficient of x term)

OpenStudy (anonymous):

thank you guys

OpenStudy (anonymous):

;)

Nnesha (nnesha):

(; o^_^o

OpenStudy (anonymous):

o^_^o